OpenStudy (anonymous):

Use a related equation that has the variable on one side. Then simplify the other side. j – 5 = 61 A. j = 56 B. j = 12.2 C. j = 66 D. j = 305

OpenStudy (anonymous):

\[j - 5 = 61=> j = 61 +5 = 6\] \[j =66\] Choice C

OpenStudy (acxbox22):

adding 5 on both sides we get j=61+5 simplifying to j=66 choice c is right

OpenStudy (anonymous):

kk thinks

OpenStudy (anonymous):

What is the value of the variable? 19b = 38 b =

OpenStudy (acxbox22):

divide by 19 on both sides and we are left with b=2

OpenStudy (anonymous):

b = 38/19 = 2

OpenStudy (anonymous):

think you

OpenStudy (anonymous):

are you shoe it is =2

OpenStudy (acxbox22):

yes

OpenStudy (anonymous):

got a 80%